# Direct Bias-Correction Term Estimation for Propensity Scores and Average Treatment Effect Estimation

### v1

@@ F.1 Preliminary
The following proposition is originally presented in van de Geer (2000) and was rephrased by Kanamori et al. (2012) in a form that is convenient for the theoretical analysis in DRE.
Lemma F.1 (Lemma 5.13 in van de Geer (2000), Proposition 1 in Kanamori et al. (2012)).
@@
and that $H_{B}(\delta,\mathcal{F}_{M},P)=O{M/\delta}^{\gamma}$. Then, we have
@@ F.2 Upper bound using the empirical-process arguments
[From Lemma 2 in Kanamori et al. (2012)] Under the conditions of Theorem 3.5, for any $0<\gamma<2$, we have
$\left|\int\left(\widehat{r}(d,\cdot)-r^{*}(d,\cdot))({\mathrm{d}}(P_{0}-P_{n})\right)\right|$ $=O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left\|\widehat{r}(d,\cdot)\right\|^{\gamma/2}_{{\mathcal{H}}}}{\sqrt{n}},\frac{\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}}{n^{2/(2+\gamma)}}\right\}\right),$
$\left|\int\left(\widehat{r}^{2}(d,\cdot)-r^{*}(d,\cdot)^{2}\right)({\mathrm{d}}(P_{0}-P_{n}))\right|$ $=O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left\|\widehat{r}(d,\cdot)\right\|^{2}_{{\mathcal{H}}}}{n^{2/(2+\gamma)}}\right\}\right),$
@@
From Proposition E.2 and $r_{0}\in{\mathcal{R}}^{\mathrm{RKHS}}$, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$
$\leq\sum_{d\in\{1,0\}}\left|\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})\right|+\left|\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\right|+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$
$=\sum_{d\in\{1,0\}}O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right)+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda),$ (4)
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$ (5)
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$ (6)
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=O_{p}(\lambda),$
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda).$
Therefore, we have $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=O_{p}(\lambda^{1/2})$ and $\|\widehat{r}\|_{{\mathcal{H}}}=O_{p}(1)$.
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right).$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{n^{1/(2+\gamma)}}\right).$
@@
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$ $=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=\sum_{d\in\{1,0\}}O_{p}\left(\left(\frac{1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}}{n^{1/(2+\gamma)}}\right)^{1-\gamma/2}\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
@@
$\|\widehat{r}\|_{{\mathcal{H}}}$ $=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{\lambda^{1/2}n^{2/(2+\gamma)}}\right)=o_{p}(1).$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
Case 6.
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right),$
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
@@
$\widehat{{\mathcal{F}}}^{f^{*},u}$ $\coloneqq\{f\in{\mathcal{F}}^{\mathrm{FNN}}\colon\frac{1}{n}\sum^{n}_{i=1}(f(X_{i})-f^{*}(X_{i}))^{2}\leq u\},$
@@
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,x)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
$\leq C\left(\|f^{*}(X)-f_{0}(D,X)\|^{2}_{2}+\|f^{*}(X)-f_{0}(X)\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}\right).$
@@
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d\mid\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
@@
Suppose that the conditions in Lemma G.1 hold. If $n\geq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, $r_{0}\geq 1/n$, and $n\geq(2eM)^{2}$, we have
${\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]\leq Cr_{0}\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log n}{n}}.$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}\leq u_{1}.$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{2}$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{2}\leq 2^{j-1}\overline{u}.$ (11)
@@
$\int-2\Big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\Big){\mathrm{d}}P_{n}$
$\ \ \ \ \ \ \leq\int-2\Big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\Big){\mathrm{d}}P_{0}+\sqrt{2}C_{1}\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{1}Mz}{3n},$
$\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ \ \ \ \leq\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\sqrt{2}C_{2}\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}.$
This is a direct consequence of Proposition D.1. Note that $r^{*}$ and $r_{0}$ are fixed, and it is enough to apply the standard law of large numbers; that is, we do not have to consider the uniform law of large numbers. However, we can still apply Proposition D.1, which is a general than the standard law of large numbers, with ignoring the Rademacher complexity part. Here, we also used
$\Big|r^{*}(d,x)-r_{0}(d,x)\Big|\leq\frac{1}{4}\Big|f^{*}(x)-f_{0}(x)\Big|,$
$\Big|r^{*}(d,x)^{2}-r_{0}(d,x)^{2}\Big|\leq\frac{1}{2}\Big|f^{*}(x)-f_{0}(x)\Big|.$
@@
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
$\leq\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{0}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
$\ \ \ \ \ +\sqrt{2}C_{1}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}+\sqrt{2}C_{2}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}\Bigg)$
@@
Additionally, we assumed that $\|\widehat{f}-f^{*}\|_{2}\leq r_{0}$ holds. Then, it holds that $\mathrm{Var}_{P_{0}}(g)\leq 4M^{2}r^{2}$.
@@
Then, from Proposition D.1, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that, for each $d\in\{1,0\}$ it holds that Then, it holds that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d\mid\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$

### v2

@@ F.1 Preliminary
We introduce the following propositions from van de Geer (2000), Kanamori et al. (2012) and Kato & Teshima (2021).
Definition F.1 (Derived function class and bracketing entropy (from Definition 4 in Kato & Teshima (2021))).
Given a real-valued function class $\mathcal{F}$}, define $\ell\circ{\mathcal{F}}:=\{\ell\circ f\colon f\in{\mathcal{F}}\}$. By extension, we define $I:\ell\circ{\mathcal{H}}\to[1,\infty)$ by $I(\ell\circ h)=I(h)$ and $\ell\circ{\mathcal{H}}_{M}:=\{\ell\circ h:h\in{\mathcal{H}}_{M}\}$. Note that, as a result, $\ell\circ{\mathcal{H}}_{M}$ coincides with $\{\ell\circ h\in\ell\circ{\mathcal{H}}:I(\ell\circ h)\leq M\}$.
Proposition F.1.
Let $\ell:{\mathbb{R}}\to{\mathbb{R}}$ be a $v$-Lipschitz continuous function. Let $H_{B}\big(\delta,{\mathcal{F}},\|\cdot\|_{L_{2}(P_{0})}\big)$ denote the bracketing entropy of $\mathcal{F}$} with respect to a distribution $P$. Then, for any distribution $P$, any $\gamma>0$, any $M\geq 1$, and any $\delta>0$, we have
$H_{B}\big(\delta,\ell\circ{\mathcal{H}},\|\cdot\|_{L_{2}(P_{0})}\big)$ $\leq\frac{(s+1)(2v)^{\gamma}}{\gamma}\left(\frac{M}{\delta}\right)^{\gamma}.$
Moreover, there exists $M>0$ such that for any $M\geq 1$ and any distribution $P$,
$\sup_{\ell\circ h\in\ell\circ{\mathcal{H}}_{M}}\|\ell\circ h-\ell\circ h^{*}\|_{L_{2}(P_{0})}$ $\leq c_{0}vM,$
$\sup_{\stackrel{{\scriptstyle\ell\circ h\in\ell\circ{\mathcal{H}}_{M}}}{{\|\ell\circ h-\ell\circ h^{*}\|_{L_{2}(P_{0})}\leq\delta}}}\|\ell\circ h-\ell\circ h^{*}\|_{\infty}$ $\leq c_{0}vM,\quad\text{for all }\delta>0.$
Proposition F.2 (Lemma 5.13 in van de Geer (2000), Proposition 1 in Kanamori et al. (2012)).
@@
and that $H_{B}(\delta,{\mathcal{F}}_{M},P)=O\left((M/\delta)^{\gamma}\right)$. Then, we have
@@
Proposition F.3.
Let $g\colon{\mathcal{K}}\to{\mathbb{R}}$ be twice continuously differentiable and strictly convex for the space $\mathcal{K}$} of $h_{0}$, and suppose that there exists $M>0$ such that
$|g^{\prime\prime}(t)|\leq M\quad\text{for all }t\in{\mathbb{R}}.$
Let $\zeta^{-1}\colon{\mathbb{R}}\to{\mathbb{R}}$ be continuously differentiable and globally Lipschitz, that is, there exists $L_{\zeta}>0$ such that
$|\zeta^{-1}(s)-\zeta^{-1}(t)|\leq L_{\zeta}|s-t|\quad\text{for all }s,t\in{\mathbb{R}}.$
Assume also that $\zeta^{-1}(0)$ is finite, and define
$a_{0}:=|\zeta^{-1}(0)|,\qquad a_{1}:=L_{\zeta},$
so that
$|\zeta^{-1}(u)|\leq a_{0}+a_{1}|u|\quad\text{for all }u\in{\mathbb{R}}.$
Let $h$ be a bounded real-valued function on the domain of $(D,X)$, and write
$\|h\|_{\infty}:=\sup_{d,x}|h(d,x)|.$
Let $L$ be a linear functional acting on bounded functions, such that for some constant $C_{L}>0$,
$|L(f)|\leq C_{L}\bigl(1+\|f\|_{\infty}\bigr)\quad\text{for all bounded }f.$
Define
$L(\zeta^{-1}\circ f)$ $=g\bigl(\zeta^{-1}\circ f(D,X)\bigr)+\partial g\bigl(\zeta^{-1}\circ f(D,X)\bigr)\,\zeta^{-1}\circ h(D,X)$
$\quad-\partial g\bigl(\zeta^{-1}\circ f(1,X)\bigr)-\partial g\bigl(\zeta^{-1}\circ f(0,X)\bigr).$
Then there exists a constant $C>0$ (depending only on $g$, $\zeta^{-1}$ and $C_{L}$) such that
$|L(\zeta^{-1}\circ f)|\leq C\bigl(1+\|f\|_{\infty}^{2}\bigr).$
@@ F.2 Upper bound using the empirical-process arguments
Under the conditions of Theorem 4.1, for any $0<\gamma<2$, we have
$d\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$=O_{p}\left(\max\left\{\frac{\|\widehat{h}_{n}-h^{*}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{h}_{n}\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{h}_{n}\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right),$
@@
From Proposition E.3 and $h_{0}\in{\mathcal{H}}^{\mathrm{RKHS}}$, we have
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{h}\|^{2}_{{\mathcal{H}}}$
$\leq\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]+\lambda\|f_{0}\|^{2}_{{\mathcal{H}}}.$
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}$
$=\O _{p}\left(\max\left\{\frac{\|\widehat{h}-h_{0}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{h}\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right)+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda),$ (2)
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\|\widehat{f}-f_{0}|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$ (3)
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$ (4)
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=O_{p}(\lambda),$
$\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda).$
Therefore, we have $\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{P_{0}}=O_{p}(\lambda^{1/2})$ and $\|\widehat{r}\|_{{\mathcal{H}}}=O_{p}(1)$.
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=O_{p}\left(\frac{\|\widehat{f}_{n}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$
$\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\|\widehat{f}_{n}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right).$
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{n^{1/(2+\gamma)}}\right).$
@@
$\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}$ $=O_{p}\left(\frac{\|\widehat{f}_{n}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=O_{p}\left(\left(\frac{1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}}{n^{1/(2+\gamma)}}\right)^{1-\gamma/2}\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
@@
$\|\widehat{f}\|_{{\mathcal{H}}}$ $=O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{2}}{\lambda^{1/2}n^{2/(2+\gamma)}}\right)=o_{p}(1).$
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{P_{0}}=O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
Case 4.
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right),$
$\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{P_{0}}=O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
@@
$\widehat{{\mathcal{F}}}^{f^{*},u}$ $\coloneqq\{f\in{\mathcal{F}}^{\mathrm{FNN}}\colon\frac{1}{n}\sum^{n}_{i=1}(f(D_{i},X_{i})-f^{*}(D_{i},X_{i}))^{2}\leq u\},$
@@
$\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\leq C\left(\|f^{*}(D,X)-f_{0}(D,X)\|^{2}_{2}+\|f^{*}(D,X)-f_{0}(D,X)\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}\right).$
@@
$\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u_{0}}\right]+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
@@
Suppose that the conditions in Lemma G.1 hold. If $n\geq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, $u_{0}\geq 1/n$, and $n\geq(2eM)^{2}$, we have
${\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u_{0}}\right]\leq Cr_{0}\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log n}{n}}.$
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}$
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}\leq u_{1}.$
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{2}$
@@
$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{2}\leq 2^{j-1}\overline{u}.$ (9)
@@
$\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\leq{\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]+\sqrt{2}C\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{1}Mz}{3n}.$
This is a direct consequence of Proposition D.1. Note that $h^{*}$ and $h_{0}$ are fixed, and it is enough to apply the standard law of large numbers; that is, we do not have to consider the uniform law of large numbers. However, we can still apply Proposition D.1, which is a general than the standard law of large numbers, with ignoring the Rademacher complexity part.
@@
$\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\leq{\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\ \ \ \ \ +\sqrt{2}C_{1}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}+\sqrt{2}C_{2}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}$
@@
Additionally, we assumed that $\|\widehat{f}-f^{*}\|_{2}\leq u_{0}$ holds. Then, it holds that $\mathrm{Var}_{P_{0}}(g)\leq 4M^{2}u^{2}_{0}$.
@@
Then, from Proposition D.1, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
$\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$